OpenStudy (anonymous):

If DH=3x-3 and FH=X+7, find the value of x for whichDEFG must be a parallelogram.

OpenStudy (anonymous):



OpenStudy (anonymous):

I got 5 as the answer

OpenStudy (mathstudent55):

What do you know about the diagonals of a parallelogram?

OpenStudy (anonymous):

not always congruent

OpenStudy (mathstudent55):

Correct, but they bisect each other. That means for this to be a parallologram, you need 3x - 3 = x + 7 2x = 10 x = 5 You are correct.